A box contains five red balls, six green balls, and nine yellow balls. Suppose you select one ball at random from the box and do not replace it. Then you randomly select a second ball.
What is the probability that both balls selected are red?

Answer by stanbon(75887)

A box contains five red balls, six green balls, and nine yellow balls. Suppose you select one ball at random from the box and do not replace it. Then you randomly select a second ball.
What is the probability that both balls selected are red?
------
P(red and red) = 5C2/20C2 = 10/190 = 10/19
============
Cheers,
Stan H.